OpenStudy (unklerhaukus):

\[ \mathcal{L} \{ erfc(\sqrt t) \}\] \[ =\mathcal{L} \{ {2\over√π}\int_{√t}^∞e^{-u^2}du \}\] \[ = {2\over√π} \int_0^∞[ \int_{√t}^∞e^{-u^2}du] e^{-st}dt\] \[ = {2\over√π} \int_0^∞[ \int_{0}^{u^2}e^{-st}dt] e^{-u^2}du\] \[ = {2\over√π} \int_0^∞ {e^{-st}\over-p} {|}^{u^2}_0 e^{-u^2}du\] \[ = {2\over√π} \int_0^∞ {1-e^{-u^{2}s}\over p} e^{-u^2}du\]

OpenStudy (unklerhaukus):

am i doing this right

OpenStudy (jamesj):

When you changed order of integration, I don't think you've got the limits right.

OpenStudy (unklerhaukus):

ok 

OpenStudy (jamesj):

\[ \int_0^\infty \int_{\sqrt{t}}^\infty ... \ du \ dt = \int_0^\infty \int_{u^2}^\infty ... dt \ du \]

OpenStudy (unklerhaukus):

isnt your right hand side that this region instead?